OpenStudy (anonymous):

Use Euler’s formula to find the missing number. Checking Answers! Vertices: 13 Edges: 26 Faces: ? 14 15 16 17<--

OpenStudy (anonymous):

No, it's 15. F+V-E=2 15+13=28 28-26=2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Your welcome.